OpenStudy (anonymous):

Trigonometry Help Prove the Identity sin 6x tan 3x = 2sin^2 3x This just isn't making too much sense to me :\

OpenStudy (anonymous):

I've tried to work the problem a bit but i just dont think its correct to me

OpenStudy (anonymous):

half angle identity is useful here

OpenStudy (anonymous):

see i don't think ive ever learned anything about half angle identity, people keep saying it but it just doent ring any bells in my head.

OpenStudy (mathstudent55):

You need a double-angle identity. sin(2x) = 2sin(x)cos(x) Use the above identity for sin 6x on the left side. Think of sin 6x as sin (2 * 3x)

OpenStudy (anonymous):

kinda weird but it works out most beautifully. i'm sure someone with the patience to use latex will explain it better than i can

OpenStudy (anonymous):

alright thanks, let me give that a try mathstudent

OpenStudy (anonymous):

erm not gettin it :\

OpenStudy (mathstudent55):

Double angle identity: \( \sin 2x = 2 \sin x \cos x\) \( \sin 6x \tan 3x ~~~~~~~~~~~~~~~~~~~~~~~~~~~~= 2 \sin^2 3x\) Think of \(\sin 6x\) as \(\sin (2 \times 3x) \) \(\sin (2 \times 3x) \tan 3x ~~~~~~~~~~~~~~~~~~~~= 2 \sin^2 3x\) \(2 \sin 3x \cos 3x \tan 3x ~~~~~~~~~~~~~~~~= 2 \sin^2 3x\) \(2 \sin 3x \cos 3x \dfrac{\sin 3x}{\cos 3x} ~~~~~~~~~~~~~~~= 2 \sin^2 3x\) \(2 \sin 3x \cancel{\cos 3x} \dfrac{\sin 3x}{\cancel{\cos 3x}~~1} ~~~~~~~~= 2 \sin^2 3x\) \(2 \sin 3x \sin 3x ~~~~~~~~~~~~~~~~~~~~~~~~~~~= 2 \sin^2 3x\) \(2 \sin^2 3x ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~= 2 \sin^2 3x\)

OpenStudy (anonymous):

you left but yeah i guess that makes sense, thank you very much

OpenStudy (mathstudent55):

You're welcome.